# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

### v3

@@ V-C3 Number of Refinement Iterations
Figure 6: Curves of average SR (left) and completeness (right) over the number of refinement iterations for the NL method on Advanced tasks.
@@ V-D Evaluator Configuration Analysis
As the precision of the evaluation feedback directly influences the quality of code refinement, we also analyzed how different evaluator configurations affect evaluation precision. The evaluation precision is defined as the ratio of correct evaluations (i.e., correctly identifying whether the trajectory aligns with the task and providing a detailed explanation if deviations are identified) to the total number of evaluations conducted. To perform this analysis, we built a dataset from each Advanced task with one correct NL trajectory observation (Correct O) and one incorrect NL trajectory observation (Incorrect O). As presented in Table III, we tested four configurations of the evaluator system prompt: (1) roles; (2) roles $+$ rules; (3) roles $+$ references; (4) roles $+$ rules $+$ references. Moreover, each configuration was evaluated three times to reduce the randomness of LLM’s output [42].
TABLE III: Evaluator Configurations
Configuration (1) (2) (3) (4)

### v4

@@ V-C2 Number of Refinement Iterations
Figure 8: Curves of SR (left) and completeness (right) over the number of refinement iterations for our method on advanced tasks.
@@ V-D Evaluator Configuration Analysis
As the accuracy of the evaluation feedback directly influences the quality of code refinement, we analyze how the evaluator’s design influences both evaluation accuracy and overall system performance. We evaluated four configurations of the evaluator’s system prompt, including (1) roles, (2) roles ++ rules, (3) roles ++ references, and (4) roles ++ rules ++ references. As specified in Section IV-D, the roles define the LLM to be an evaluator and serve as the essential component to ensure the LLM will perform the evaluation task. Thus, we include the roles in each configuration of the evaluator. For each configuration, we instantiate the evaluator’s system prompt by including only the indicated components. The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table IV.
TABLE V: Performance over Evaluator Configurations
(1) (2) (3) (4)